# Corks, exotic 4-manifolds and genus functions

### v3

@@ 9. Exotic 4-manifolds and genus functions
We first show that any $4$-dimensional $2$-handlebody can be modified into infinite order generalized corks. The theorem below is a refinement of Theorem 4.1.
@@
For any $4$-dimensional $2$-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ topologically resembles $X$, admits an embedding into $X$, and is homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$. Furthermore, in the case where $H_{2}(X)$ is a non-trivial group, there exist arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that $X_{i}$’s have the properties of $X^{\prime}$ and are pairwise exotic, and that the genus functions of $X_{i}$’s are pairwise algebraically inequivalent for any orientations.
@@
Let $X$ be a 4-dimensional 2-handlebody. By Proposition 3.3, we have a 4-manifold $Y$ admitting a Stein structure such that $Y$ topologically resembles $X$, admits an embedding into $X$, and is homeomorphic to a 4-manifold admitting a deformation retraction onto $X$. Therefore, by Proposition 4.2, we obtain an infinite order generalized cork $(X^{\prime},f)$ having the required properties.
Now we assume $H_{2}(X)$ is non-trivial. Then, by Theorem 3.4, we obtain arbitrarily many pairwise exotic 4-manifolds $Y_{i}$’s such that each $Y_{i}$ have the above properties of $Y$ and that the genus functions of $Y_{i}$’s are pairwise algebraically inequivalent for any orientations. Therefore, by the construction in the proof of Proposition 4.2, we obtain arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that each $X_{i}$ have the above properties of $X^{\prime}$ and that $X_{i}$’s are pairwise homeomorphic. Since each $X_{i}$ is obtained by attaching an invertible homology cobordism to $Y_{i}$, the genus function of each $X_{i}$ is algebraically equivalent to that of $Y_{i}$ due to Corollary 7.5, and thus $X_{i}$’s are pairwise exotic. This completes the proof. ∎
@@
Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks with $\partial C=M$, and assume that each $f\in A$ extends to a self-homeomorphism of $C$ that preserves the genus function. Then there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic $4$-manifolds topologically resembling $C$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $C$. Furthermore, they admit embeddings into $C$ and deformation retractions onto $C$.
@@
By Corollary 3.10, we have a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic 4-manifolds having all the desired properties except those on genus functions. We note that $C_{\mathrm{id}_{M}}$ is obtained from $C$ by attaching an invertible cobordism $\mathbb{P}$, and each $C_{f}$ is obtained from $C_{\mathrm{id}_{M}}$ by removing $C$ and gluing it back using $f$. By Corollary 7.5, the genus function of each $C_{f}$ is algebraically equivalent to that of $C$, and this equivalence is induced by the inclusion $C\hookrightarrow C_{f}$. Since each $f\in A$ preserves the orientation and extends to a self-homeomorphism of $C$ that preserves the genus function, this self-homeomorphism of $C$ extends to an orientation-preserving homeomorphism $C_{\mathrm{id}_{M}}\to C_{f}$ that preserves the genus functions. ∎
Now we will prove results on exotic 4-manifolds given in Section 1. The theorem below is a refinement of Theorem 1.3.
@@
For any $4$-dimensional $2$-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent. Furthermore, they admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
@@
Suppose that a $4$-manifold $X$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$. Furthermore, they admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
@@
The corollary below is a slight refinement of Corollary 1.6.
@@
$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((((with boundary)))) such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
$(2)$ If $X$ is the total space of a $D^{2}$-bundle over a closed surface, then there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$.
@@
(1) Since $D^{4}$ is a Stein manifold, its blow-ups $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is a symplectic $4$-manifold with weakly convex boundary. Also, the genus function of $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is equivalent to that of $n\overline{\mathbb{C}P^{2}}$ due to Lemma 7.2. The claim (1) thus follows from Theorem 9.5.
(2) Let $X$ denote the total space of $D^{2}$-bundle over a closed surface of genus $g$ with Euler number $n$. Then $X$ admits a Stein structure for $n\leq 2g-2$ (see [38, Exercises 11.2.5.(a)]). Hence the claim (2) follows from Theorem 9.5 in the case where $n\leq 2g-2$. In the case where $n\geq 2g-1$ and $g\geq 1$, the claim (2) follows from the case where $n\leq 2g-2$ by reversing the orientation. The same argument also shows the claim (2) in the case where $n\geq 2$ and $g=0$. The remaining cases are $(g,n)=(0,0),(0,\pm 1)$. In the case where $(g,n)=(0,\pm 1)$, the claim (2) follows from (1). In the case where $(g,n)=(0,0)$, the 4-manifold $X$ is diffeomorphic to $S^{2}\times D^{2}$ and hence admits an embedding into $D^{4}$. The claim (2) thus follows from Theorem 9.5 in this case. This completes the proof. ∎
@@
There exist infinitely many pairwise exotic simply connected $4$-manifolds such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is orientedly equivalent to those of another infinitely many pairwise exotic $4$-manifolds.
@@
Let $\{X_{i}\}_{i\in\mathbb{Z}}$ be an infinite family of pairwise exotic 4-manifolds admitting Stein structures such that the genus functions of the members are algebraically inequivalent for any orientations. Such examples were obtained by Akbulut and the author in [12] and the author in [68]. By Proposition 4.2, we obtain an infinite family $\{(Y_{i},f_{i})\}_{i\in\mathbb{Z}}$ of infinite order generalized corks such that each $Y_{i}$ is obtained from $X_{i}$ by attaching the same invertible homology cobordism, and that $Y_{i}$’s are pairwise homeomorphic. By Corollary 7.5, the genus function of each $Y_{i}$ is algebraically equivalent to that of $X_{i}$. Due to the proof of Proposition 9.3, by attaching another fixed invertible homology cobordism to each $Y_{i}$ and twisting $Y_{i}$ in the resulting 4-manifold $X^{\prime}_{i,0}$ using each $f_{i}^{j}$, we obtain an infinite family $\{X^{\prime}_{i,j}\}_{(i,j)\in\mathbb{Z}\times\mathbb{Z}}$ of pairwise exotic simply connected 4-manifolds. By Proposition 9.3, for each fixed $i$, the genus functions of $X^{\prime}_{i,j}$’s are pairwise equivalent, and furthermore for each fixed $j$, the genus functions of $X^{\prime}_{i,j}$’s are pairwise algebraically inequivalent for any orientations. ∎
@@
Here we construct similar examples for much more homeomorphism types, by relaxing the condition “infinitely many” to “arbitrarily many”. The theorem below is a refinement of Theorem 1.8.
@@
For any $4$-dimensional $2$-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is orientedly equivalent to those of infinitely many pairwise exotic $4$-manifolds. Furthermore, all of them admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
@@
As an application of our results on genus functions, we give exotic 4-manifolds which are stably exotic under boundary sums with arbitrary Stein manifolds having non-degenerate intersection forms. The theorem below is a restatement of Theorem 1.11.
@@
For any positive integer $n$, there exist pairwise exotic simply connected $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that, for any pairwise orientedly homeomorphic $4$-manifolds $Y_{1},Y_{2},\dots,Y_{n}$ having non-degenerate intersection forms and admitting Stein structures, the boundary sums $X_{1}\natural Y_{1}$, $X_{2}\natural Y_{2}$, $\dots$, $X_{n}\natural Y_{n}$ are pairwise exotic.
@@
Let $X$ be a 4-dimensional 2-handlebody with non-trivial second homology group such that its intersection form is represented by a zero-matrix, and let $Y_{1},Y_{2},\dots,Y_{n}$ be 4-manifolds having non-degenerate intersection forms and admitting Stein structures. By Theorem 3.4, we have pairwise exotic $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that they topologically resemble $X$ and that their genus functions are pairwise algebraically inequivalent for any orientations. Since each $X_{i}$ and $Y_{i}$ admit Stein structures, we can apply the argument on estimating genus functions in the proof of Theorem 3.4, and thus we see that the restrictions of the genus functions of $X_{i}\natural Y_{i}$’s to $H_{2}(X_{i})$’s remain pairwise algebraically inequivalent. Note that the second homology group of any 2-handlebody is torsion-free. It thus follows from Proposition 5.2 that the genus functions of $X_{i}\natural Y_{i}$’s are pairwise inequivalent. ∎
As seen from the above proof, we can realize various homeomorphism types as those of $X_{i}$’s.
@@ 10. Exotically knotted submanifolds
As an application of our exotic 4-manifolds in the last section, we modify a submanifold of a 4-manifold with codimension less than two into infinitely many pairwise exotically knotted submanifolds.
For a codimension-00 submanifold $X$ of a 4-manifold $Z$, the closure of $Z-X$ will be called the exterior of $X$ and denoted by $E(X)$. To state our results shortly, we use the following terminologies.
@@
$(1)$ The pair $X_{1},X_{2}$ will be called exotically knotted if they have the following properties.
@@
$(2)$ The pair $X_{1},X_{2}$ will be called exotically embedded if they have the following properties.
@@
We produce exotically knotted submanifolds of codimension-0. The theorem below is a refinement of Theorem 1.12.
@@
Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$. These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries. Suppose the exterior of $X$ is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-00 submanifolds topologically resembling $X$ such that each of these $4$-manifolds contains $X$ as a submanifold and deformation retracts to $X$.
To prove this theorem, we give a method for producing exotically knotted submanifolds by using generalized corks obtained from contractible corks. We use the following setting only for the lemma and the proposition below. Let $\{(C_{0},f)\}_{f\in A}$ be a family of contractible corks with connected boundary having a fixed embedded 3-ball $B_{0}$ in $\partial C_{0}$ such that every $f\in A$ is identity on a neighborhood of the 3-ball $B_{0}$. Let $Y$ be a 4-manifold with $\partial Y=M$. Take the boundary sum $(I\times M)\natural C_{0}$ for the boundary component $1\times M$ using the 3-ball $B_{0}$ of $\partial C_{0}$. This manifold $P$ is a homology cobordism from $M$ to $M\#\partial C_{0}$. We set $M^{\prime}=M\#\partial C_{0}$. Abusing the notation, let $f$ denote the self-diffeomorphism of $M^{\prime}$ that extends the self-diffeomorphism $f$ of $\partial C_{0}$ as the identity on the rest of $M^{\prime}$. We denote the glued 4-manifold $Y\cup_{\mathrm{id}_{M}}P$ by $C$, and assume $\{(C,f)\}_{f\in A}$ is a family of pairwise inequivalent generalized corks. We furthermore assume that, for each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, the twisted double $C_{0}\cup_{f}\overline{C_{0}}$ of $C_{0}$ is diffeomorphic to $S^{4}$.
We identify the double of $I\times M$ with $[0,2]\times M$. Then, for each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, the twisted double $P\cup_{f}\overline{P}$ can be regarded as the connected sum of $[0,2]\times M$ and the twisted double $C_{0}\cup_{f}\overline{C_{0}}$. Since $f$ is identity on a neighborhood of the 3-ball $B_{0}$ in $\partial C_{0}$, it follows that the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ has an embedded 4-ball $D_{P}$ such that every twisted double $P\cup_{f}\overline{P}$ is obtained from $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{P}$ by attaching a “twisted partial double” of $C_{0}$, which is obtained from $C_{0}$ by attaching $\overline{C_{0}}$ using the restriction of $f$ to the exterior of $B_{0}$. We here observe the following lemma.
@@
$(2)$ For each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, there exists a diffeomorphism $P\cup_{f}\overline{P}\to I\times M$ such that it fixes the boundary $M\sqcup\overline{M}$ pointwise and that its restriction to the submanifold $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{P}$ is independent of $f$.
@@
The claim $(1)$ is straightforward from the construction of $P$. Since each twisted double $C_{0}\cup_{f}\overline{C_{0}}$ is diffeomorphic to $S^{4}$, the disk theorem shows that the twisted partial double of $C_{0}$ is diffeomorphic to $D^{4}$. The claim (2) thus follows from the decomposition of the twisted double of $P$ in the paragraph preceding this lemma. ∎
@@
Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$. These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries. If the exterior $E(X)$ is diffeomorphic to $Y$, then $Z$ admits a family $\{X_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}}$ of pairwise exotically knotted codimension-00 submanifolds topologically resembling $X$ such that each $X_{f}$ contains $X$ as a submanifold and admits a deformation retraction onto $X$.

### v5

@@ 9. Exotic 4-manifolds and genus functions
We first show that any $4$-dimensional 22-handlebody can be modified into infinite order generalized corks. Specifically, we prove the following theorem which is a refinement of Theorem 4.1.
@@
For any $4$-dimensional 22-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ is HIHC-equivalent to $X$, admits an embedding into $X$, and is homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$. Moreover, in the case where $H_{2}(X)\neq 0$, there exist arbitrarily many generalized corks $(X_{i},f_{i})$ of infinite order with the properties of $(X^{\prime},f)$ such that the $4$-manifolds $X_{i}$ are pairwise exotic. Furthermore, their genus functions are pairwise algebraically inequivalent for any orientations.
@@
Let $X$ be a 4-dimensional 2-handlebody. By Proposition 3.6, we have a 4-manifold $Y$ admitting a Stein structure such that $Y$ is HIHC-equivalent to $X$, admits an embedding into $X$, and is homeomorphic to a 4-manifold that deformation retracts onto an embedded copy of $X$. Therefore, by Proposition 4.2, we obtain an infinite order generalized cork $(X^{\prime},f)$ having the required properties.
Now we assume $H_{2}(X)\neq 0$. Then, by Theorem 3.7, we obtain arbitrarily many pairwise exotic 4-manifolds $Y_{i}$ such that each $Y_{i}$ have the above properties of $Y$ and that the genus functions of $Y_{i}$ are pairwise algebraically inequivalent for any orientations. Therefore, by the construction in the proof of Proposition 4.2, we obtain arbitrarily many generalized corks $(X_{i},f_{i})$ of infinite order such that the 4-manifolds $X_{i}$ have the above properties of $X^{\prime}$ and are pairwise homeomorphic. Since each $X_{i}$ is obtained by attaching an invertible homology cobordism to $Y_{i}$, the genus function of each $X_{i}$ is algebraically equivalent to that of $Y_{i}$ due to Corollary 7.5, and thus the 4-manifolds $X_{i}$ are pairwise exotic. This completes the proof. ∎
@@
Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks, and assume each $f\in A$ extends to a self-homeomorphism of $C$ that preserves the genus function. Then, there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ of pairwise exotic $4$-manifolds HIHC-equivalent to $C$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $C$. Furthermore, they admit deformation retractions onto an embedded copy of $C$ and embeddings into $C$.
@@
For each $f\in A\sqcup\{\mathrm{id}_{\partial C}\}$, let $C_{f}$ denote the 4-manifold $C\cup_{f}\mathbb{P}$ given in the proof of Corollary 3.13. By Corollary 3.13 and Lemma 3.11, the family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ has all the desired properties possibly except those on genus functions. By Corollary 7.5, the genus function of each $C_{f}$ is algebraically equivalent to that of $C$, and this equivalence is induced by the inclusion $C\hookrightarrow C_{f}$. Since each $f\in A$ preserves the orientation and extends to a self-homeomorphism of $C$ that preserves the genus function, this self-homeomorphism of $C$ extends to an orientation-preserving homeomorphism $C_{\mathrm{id}_{\partial C}}\to C_{f}$ that preserves the genus functions. Therefore, the family has the desired properties on genus functions as well. ∎
Now we are ready to produce various infinite families of exotic 4-manifolds with pairwise equivalent genus functions. We here prove the following theorem which is a refinement of Theorem 1.3.
@@
For any $4$-dimensional 22-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that their genus functions are pairwise orientedly equivalent. Furthermore, these $4$-manifolds admit embeddings into $X$ and are homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$.
@@
For any $4$-manifold $X$ admitting an embedding into a symplectic $4$-manifold with weakly convex boundary, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$. Furthermore, these $4$-manifolds admit deformation retractions onto embedded copies of $X$ and embeddings into $X$.
@@
The above theorem implies the following corollary, which is a slight refinement of Corollary 1.6.
@@
$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((with boundary)) such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
$(2)$ For any $D^{2}$-bundle over a closed surface, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to the total space $X$ of the bundle such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$.
@@
(1) Since $D^{4}$ admits a Stein structure, its blow-ups $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is a symplectic $4$-manifold with weakly convex boundary. Also, the genus function of $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is equivalent to that of $n\overline{\mathbb{C}P^{2}}$ due to Lemma 7.2. The claim (1) thus follows from Theorem 9.5.
(2) Let $X$ denote the total space of the $D^{2}$-bundle over a closed surface of genus $g$ with Euler number $n$. Then $X$ admits a Stein structure for $n\leq 2g-2$ (see [39, Exercises 11.2.5.(a)]). Hence the claim (2) follows from Theorem 9.5 in the case where $n\leq 2g-2$. In the case where $n\geq 2-2g$, the claim (2) follows from the case where $n\leq 2g-2$ by reversing the orientation. The remaining cases are $(g,n)=(0,0),(0,\pm 1)$. In the case where $(g,n)=(0,\pm 1)$, the claim (2) follows from (1). In the case where $(g,n)=(0,0)$, the 4-manifold $X$ is diffeomorphic to $S^{2}\times D^{2}$ and hence admits an embedding into $D^{4}$. The claim (2) thus follows from Theorem 9.5 in this case. This completes the proof. ∎
@@
There exist infinitely many pairwise exotic simply connected $4$-manifolds whose genus functions are pairwise algebraically inequivalent for any orientations such that, for each $4$-manifold $X$ among them, there exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X$ whose genus functions are all orientedly equivalent to that of $X$.
@@
Let $\{X_{i}\}_{i\in\mathbb{Z}}$ be an infinite family of pairwise exotic 4-manifolds admitting Stein structures such that the genus functions of the members are algebraically inequivalent for any orientations. Such examples were obtained by Akbulut and the author in [12] and the author in [70]. By the construction in the proof of Proposition 4.2, we obtain an infinite family $\{(Y_{i},f_{i})\}_{i\in\mathbb{Z}}$ of infinite order generalized corks such that each $Y_{i}$ is obtained from $X_{i}$ by attaching the same invertible homology cobordism, and that the 4-manifolds $Y_{i}$ are pairwise homeomorphic. By Corollary 7.5, the genus function of each $Y_{i}$ is algebraically equivalent to that of $X_{i}$. Due to the proof of Proposition 9.3, by attaching another fixed invertible homology cobordism to each $Y_{i}$ and twisting $Y_{i}$ in the resulting 4-manifold $X^{\prime}_{i,0}$ using each $f_{i}^{j}$, we obtain an infinite family $\{X^{\prime}_{i,j}\}_{(i,j)\in\mathbb{Z}\times\mathbb{Z}}$ of pairwise exotic simply connected 4-manifolds. By Proposition 9.3, for each fixed $i$, the genus functions of the 4-manifolds $X^{\prime}_{i,j}$ are pairwise equivalent, and furthermore for each fixed $j$, the genus functions of $X^{\prime}_{i,j}$ are pairwise algebraically inequivalent for any orientations. ∎
@@
Here we construct similar examples for much more homeomorphism types, by relaxing the condition “infinitely many” to “arbitrarily many”. Specifically, we prove the following theorem which is a refinement of Theorem 1.8.
@@
For any $4$-dimensional 22-handlebody $X$ with $H_{2}(X)\neq 0$, there exist arbitrarily many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ whose genus functions are pairwise algebraically inequivalent for any orientations, such that each $4$-manifold $X^{\prime}$ among them satisfies the following.
  • There exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X^{\prime}$ whose genus functions are all orientedly equivalent to that of $X^{\prime}$.
  • $X^{\prime}$ and the above infinitely many $4$-manifolds admit embeddings into $X$ and are homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$.
@@
As an application of our results and Theorem 3.7, we give exotic 4-manifolds which are stably exotic under boundary sums with arbitrary Stein manifolds having non-degenerate intersection forms. Specifically, we prove the following theorem which is a refinement of Theorem 1.11.
@@
Let $X$ be a $4$-dimensional 22-handlebody with $H_{2}(X)\neq 0$ whose intersection form is represented by a zero-matrix. Then for any positive integer $n$, there exist pairwise exotic $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ HIHC-equivalent to $X$ and admitting Stein structures such that, for any pairwise orientedly homeomorphic $4$-manifolds $Y_{1},Y_{2},\dots,Y_{n}$ having non-degenerate intersection forms and admitting Stein structures, the boundary sums $X_{1}\natural Y_{1}$, $X_{2}\natural Y_{2}$, $\dots$, $X_{n}\natural Y_{n}$ are pairwise exotic.
@@
By Theorem 3.7, we have pairwise exotic $4$-manifolds $X_{1},\dots,X_{n}$ HIHC-equivalent to $X$ and admitting Stein structures such that their genus functions are pairwise algebraically inequivalent for any orientations. Let $Y_{1},\dots,Y_{n}$ be pairwise orientedly homeomorphic 4-manifolds having non-degenerate intersection forms and admitting Stein structures. Since each $X_{i}$ and $Y_{i}$ admit Stein structures, we can apply the argument on the genus functions of the 4-manifolds $X_{i}$ in the proof of Theorem 3.7 to the restrictions of those of the 4-manifolds $X_{i}\natural Y_{i}$ to $H_{2}(X_{i})$ in the exactly same way. We thus see that the restrictions of the genus functions of $X_{i}\natural Y_{i}$ to $H_{2}(X_{i})$ remain pairwise algebraically inequivalent for any orientations. Since the second homology group of any 2-handlebody is torsion-free, it follows from Proposition 5.2 that the genus functions of $X_{i}\natural Y_{i}$ are pairwise algebraically inequivalent for any orientations. Therefore, the 4-manifolds $X_{i}\natural Y_{i}$ are pairwise exotic. ∎
There are rich examples of 4-manifolds satisfying the above assumption of $X$, hence we can realize various homeomorphism types including non-simply connected ones as those of the above $X_{1},\dots,X_{n}$.
@@ 10. Exotically knotted submanifolds
As an application of our exotic 4-manifolds in the last section, we modify a submanifold of a 4-manifold with codimension at most one into infinitely many pairwise exotically knotted submanifolds. Note that we follow Convention 1.1.
For a codimension-0 submanifold $X$ of a 4-manifold $Z$, the closure of ZXZ-X is called the exterior of $X$ and denoted by $E(X)$. To state our results shortly, we use the following terminologies.
@@
$(1)$ The pair $X_{1},X_{2}$ is called exotically knotted if they have the following properties.
@@
$(2)$ The pair $X_{1},X_{2}$ is called exotically embedded if they have the following properties.
@@
In the above definition, we require the conditions on orientations and the boundary so that our examples have stronger properties, but our results are still new without these extra conditions. Also, our results are new as exotically knotted submanifolds and also as exotically embedded submanifolds.
We first produce exotically knotted submanifolds of codimension-0. Specifically, we prove the following theorem which is a refinement of Theorem 1.12.
@@
Let $X$ be a codimension-0 submanifold of a $4$-manifold $Z$. These manifolds may be non-orientable and may have ((possibly disconnected)) boundaries. Suppose the exterior of $X$ either is a ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary, and assume the exterior has connected boundary. Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-0 submanifolds HIHC-equivalent to $X$. Furthermore, each of these submanifolds contains $X$ as a submanifold and admits a deformation retraction onto $X$.
To prove this theorem, we give a method for producing exotically knotted submanifolds by using generalized corks obtained from contractible corks. We use the following setting only for the forthcoming Lemma 10.3 and Proposition 10.4. Let $\{(C_{0},f)\}_{f\in A}$ be a family of pairwise inequivalent contractible corks having a fixed embedded 3-ball $B_{0}$ in $\partial C_{0}$ such that every $f\in A$ is identity on a neighborhood of $B_{0}$. Let $Y$ be a 4-manifold with non-empty boundary $M$. Take the boundary sum $(I\times M)\natural C_{0}$ for the boundary component $1\times M$ using the 3-ball $B_{0}$ of $\partial C_{0}$. This manifold $P$ is a homology cobordism from $M$ to $M\#\partial C_{0}$. We set $M^{\prime}=M\#\partial C_{0}$. For each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$, let $\widetilde{f}$ denote the self-diffeomorphism of $M^{\prime}$ that extends the self-diffeomorphism $f$ of $\partial C_{0}$ as the identity on the rest of $M^{\prime}$. We denote the glued 4-manifold $Y\cup_{\mathrm{id}_{M}}P$ by $C$, and assume that $\{(C,\widetilde{f})\}_{f\in{A}}$ is a family of pairwise inequivalent generalized corks. We furthermore assume that, for each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$, the twisted double $C_{0}\cup_{{f}}\overline{C_{0}}$ of $C_{0}$ is diffeomorphic to $S^{4}$.
Let $D_{C_{0}}$ be the “partial double” $C_{0}\cup_{\mathrm{id}_{\partial C_{0}-\mathrm{int}\,B_{0}}}\overline{C_{0}}$ of $C_{0}$, which is obtained from $C_{0}$ by attaching $\overline{C_{0}}$ using the identity map on the exterior of $B_{0}$. Clearly, $D_{C_{0}}$ is a submanifold of the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$. Due to the assumption on the double of $C_{0}$, the disk theorem implies that $D_{C_{0}}$ is diffeomorphic to the 4-ball. Since each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$ is identity on a neighborhood of the 3-ball $B_{0}$ in $\partial C_{0}$, it follows that the twisted double $P\cup_{\widetilde{f}}\overline{P}$ is obtained from $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{C_{0}}$ by attaching the “twisted partial double” $C_{0}\cup_{f|_{\partial C_{0}-\mathrm{int}\,B_{0}}}\overline{C_{0}}$ of $C_{0}$, which is obtained from $C_{0}$ by attaching $\overline{C_{0}}$ using the restriction of $f$ to the exterior of $B_{0}$. We note that each twisted double $P\cup_{\widetilde{f}}\overline{P}$ is diffeomorphic to the connected sum of $I\times M$ and the twisted double $C_{0}\cup_{f}\overline{C_{0}}$.
@@
$(2)$ For each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$, there exists a diffeomorphism $P\cup_{\widetilde{f}}\overline{P}\to I\times M$ such that it fixes the boundary $M\sqcup\overline{M}$ pointwise and that its restriction to the submanifold $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{C_{0}}$ is independent of $f$.
@@
The claim $(1)$ is straightforward from the construction of $P$. Since each twisted double $C_{0}\cup_{f}\overline{C_{0}}$ is diffeomorphic to $S^{4}$, the disk theorem shows that the twisted partial double of $C_{0}$ is diffeomorphic to $D^{4}$. Thus, every diffeomorphism between the boundary of the twisted partial double and $\partial D^{4}$ extends to a diffeomorphism between their interiors. Therefore, the claim (2) follows from the aforementioned decomposition of the twisted double of $P$. ∎
The following proposition gives exotically knotted codimension-0 submanifolds.
@@
Let $X$ be a codimension-0 submanifold of a $4$-manifold $Z$. These manifolds may be non-orientable and may have ((possibly disconnected)) boundaries. If the exterior $E(X)$ is diffeomorphic to the $4$-manifold $Y$, then $Z$ admits a family $\{X_{\widetilde{f}}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}}$ of pairwise exotically knotted and exotically embedded codimension-0 submanifolds HIHC-equivalent to $X$. Furthermore, each $X_{\widetilde{f}}$ contains $X$ as a submanifold and admits a deformation retraction onto $X$.